Jack is eight years older than his sister. In three years he will be twice as old as she is, how old are they now?
Answer by TimothyLamb(4379)

j = s + 8
j + 3 = 2(s + 3)
---
j + 3 = 2(s + 3)
s + 8 + 3 = 2s + 6
---
s = 5
j = 13
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php